OpenStudy (anonymous):

find the gcf of the following 46a 50a^2

OpenStudy (michele_laino):

hint: we have the subsequent factorizations: \[\Large \begin{gathered} 46a = 2 \times 23 \times a \hfill \\ 50{a^2} = 2 \times {5^2} \times {a^2} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

now, the great common factor is given by the product of common factors to those factorization, each with the littlest exponent